How much water must a landscaper add to a 10% solution of selective herbicide to create 5 gallons of 4% solution? How much of the 10% solution did the landsccaper have before he added the water?
Found 2 solutions by stanbon, richwmiller:
Answer by stanbon(75887)

How much water must a landscaper add to a 10% solution of selective herbicide to create 5 gallons of 4% solution?
--------------
Equation:
water + water = water
x + 0.90(5-x) = 0.96*5
Multiply thru by 100 and solve for "x":
100x + 90*5 - 90x = 96*5
10x = 6*5
x = 3 gallons (amt. of water needed for the mixture)
==========================
How much of the 10% solution did the landscaper have before he added the water?
What does that mean?
==========================
Cheers,
Stan H.

Answer by richwmiller(17219)

We can work this from the view of the herbicide or from the view of the water.
.10h=5*.04
.10h=.20
h=2
start with 2 gallons of herbicide and add 3 gallons of water.
THe question "How much of the 10% solution did the landscaper have before he added the water?" is poorly worded.
We don't know maybe he had just 2 gallons maybe he had a warehouse of it.
But he only used 2 gallons of the 10% herbicide.
from the view of the water
.90h+(5-h)=.96*5
.9h+5-h=4.8
.9h-h=-.2
-.1h=-.2
h=2 gallons of herbicide
water=5-2=3 so 3 gallons of water